OpenStudy (waleed_imtiaz):

prove that if n is any odd integer then 8/n^2-1 by mathematical induction?

OpenStudy (anonymous):

Just replace n with 2m+1

OpenStudy (anonymous):

Well, I'd start with that at least.

OpenStudy (anonymous):

You could also try substituting n with n+1

OpenStudy (waleed_imtiaz):

can u please give me the full solution step by step...

OpenStudy (anonymous):

Wait I don't the question is complete. You just said "prove that if n is an odd integer then 8/9^2-1"...What 8/9^2-1?

OpenStudy (waleed_imtiaz):

i said 8|n^2-1 .... prove by mathematical induction that for all values of n(odd integers) , the function is divisible by 8...

OpenStudy (anonymous):

Base case n = 1. 8|n^2-1 = 8|0 True Assume 8|n^2-1 8|(n+2)^2-1 = 8| n^2+ 4n+3 = 8|n^2-1+4n+4 = 8|n^2-1+4(n+1) which is True since n^2-1 divides 8 and 4(n+1) divides 8 since n+1 is even and 4/8(n+1)=4/8*2*n=n

OpenStudy (sirm3d):

an odd number n can be represented by either (a) 4m+1, or (b) 4m+3 use mathematical induction on each case